OpenStudy (anonymous):

Simplify. (–12x) ÷ (–2) A.12x B.1/12x C.3x D.6x

OpenStudy (michele_laino):

hint: we have this: \[\Large \frac{{ - 12x}}{{ - 2}} = \left( {\frac{{ - 12}}{{ - 2}}} \right)x = ...?\]

OpenStudy (anonymous):

D

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

YEAHHHH THANKS :D . @Michele_Laino

OpenStudy (michele_laino):

:)